A team of scientists found that there were 4 oak trees for every 10 pine trees. How many oak trees were there if they counted 36 more pine than oak?
Answer by ewatrrr(24785)


Hi
36 more pine than oak
Let x and (x+36) represent the number of oak and pine respectively
Question states***
4/10 = x/(x+36)
10x = 4(x+36)
6x = 4*36
x = 4*36/6
x = 24, the number of oaks. 60 pine. (24-36)
CHECKING our Answer***
24/60 = 4/10